Fermi National Laboratory
 

Accelerator Update for April 2, 2007 - April 4, 2007

 

Monday April 2

The day shift began with the Tevatron (TeV) in a wet squeeze, with Operations monitoring the Recycler stash (335.7E10), with the Antiproton Source (Pbar) stacking (50.6E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

At 9:44 AM, the MI beam permit tripped off for the first of three times in 15 minutes due to bad beam positions. It reset.

Operations put the TeV into shot setup at 10:08 AM.

A Pelletron expert began an Electron Cooling orbit study at 11:49 AM.

Operations established store 5317 at 11:57 AM with an initial luminosity of 243.45E30.

FESS technicians began their pond water strainer cleaning at 12:16 PM. The work went on for most of the afternoon.

Operations reported at 12:40 PM that Preaccelerator experts were switching from the H- Source to the I- Source. RF experts were bypassing a Booster RF station (BRF18).

A Pbar expert began a 30-minute stacktail measurement study at 2:02 PM.

A MI expert began a slipstacking study at 2:09 PM.

Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.

At 4:07 PM, Operations reported that the MI beam permit had been tripping off due to the MI-8 collimator all afternoon. Operators investigated, found that they were not in the correct position, and contacted a collimator expert.

At 6:28 PM, Operations reported the first of many MI permit trips due to the MI-8 collimators over the next 30 minutes. An expert commented that the problem was related to the momentum spread of the beam in the MI-8 transfer line.

At 10:49 PM, Operations reported that the vacuum gauge and the arc supply for the I- Source was tripping off. Operators lowered the extractor voltage by 1kV.

At 11:21 PM, Operations reported the second emittance blowup in the Recycler. Experts were investigating.

Operations reported at 11:58 PM that the Recycler had stashed twice during the evening shift.

Tuesday April 3

The midnight shift began with Operations monitoring store 5317 (73.79E30) and the stash (165.4E10), with Pbar stacking (69.0E10), and with NuMI and MTest taking beam.

MTest reported at 12:23 AM that they were finished with the beam for the night. This was the end of the run for T950.

A Recycler expert reported at 2:23 AM that the emittance problems were due to the new damper system they had recently switched to. He switched back to the old damper system.

Operations terminated store 5317 at 6:33 AM.

Operations had Pbar, enclosure B, and MI-65 ready for access by 6:47 AM. Booster was ready for access by 7:07 AM.

By 7:30 AM, most of the accelerator complex, except for MI/Recycler, was ready for access.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

Booster came out of access at 10:52 AM.

TeV Transfer Hall and F sector were out of access by 11:02 AM.

Cryo system technicians at CHL reported at 11:21 AM that they were starting the nitrogen plant back up.

NuMI was out of access by 11:55 AM.

At 12:34 PM, Preaccelerator experts reported that the dome temperature problems were due to old equipment that needed replacing. The arcing, cathode temperature, and extractor voltage problems were all due to extractor sparking and that was caused by the vacuum system. The slow recovery after a spark was due to the vacuum system thinking there was a leak and shutting off the extractor and arc supplies; the supplies turned on slowly. The experts made repairs.

NuMI resumed taking beam at 1:16 PM.

Pbar resumed stacking at 1:24 PM.

Pbar experts began making stacktail-cooling measurements at 1:29 PM.

At 1:59 PM, Operations reported that two Pbar devices (D:PMAGV and D:EKIK) were not reading back. Operators contacted experts.

At 2:23 PM, the MI permit tripped off for the first of five times in about 90 minutes. It reset.

At 3:01 PM, Operations reported that the Pbar tuning program didn’t begin properly. An expert investigated.

A Pbar expert began a stacking study at 7:04 PM.

Operators completed a search and secure of the D0 collision hall at 7:25 PM.

At 10:16 PM, Operations reported that a problem with a TeV power supply (RPSB1) water flow bit that would not reset. Operations called in an expert.

Operations reported at 11:57 PM that the Recycler had stashed once during the evening shift.

Wednesday April 4

The midnight shift began with Operations monitoring the stash (312.6E10), with Pbar stacking (51.0E10), and with NuMI taking beam.

The power supply expert reported at 1:07 AM that he found a connector that had fallen off a CAMAC card and that was why the bit would not reset for RPSB1.

Operations put the TeV through a dry squeeze at 1:09 AM.

At 1:47 AM, the MI permit tripped off for the first of two times in about 30 minutes. It reset okay.

Operations put the TeV through a 4x0 wet squeeze.

Operations put the TeV into shot setup at 3:33 AM.

At 4:21 AM, the TeV aborted due to beam positions at D0. Only protons were in the machine at the time and there was no quench. Investigations found no explanation for the abort.

Operations put the TeV into shot setup at 5:01 AM.

At 6:04 AM, BRF14 tripped off for the first off seven times in 9 minutes. Operators investigated and noticed the trips were due to a Pbar timeline. Operations switched to a normal stacking/NuMI timeline and the trips stopped.

At 6:21 AM, the TeV suffered a three-house quench (sectors A1, D1, and D0), which occurred after the antiprotons had been loaded and Operations had started squeezing. Experts investigated.

Operations reported at 7:51 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Wednesday

The plans for today are to recover the TeV and then establish a small store. The main priority will be stacking and stashing. The Run Coordinator will allow experts to conduct a Pbar stacktail measurement and a MI nondestructive collimator study today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/06/2007    email Fermilab